# PDMLP: Patch-based Decomposed MLP for Long-Term Time Series Forecasting

### v2

@@ 3.3 MLP layer
Methods PDMLP iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear SCINet FEDformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.316 0.357 0.334 0.368 0.355 0.376 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372 0.418 0.438 0.379 0.419
192 0.357 0.378 0.377 0.391 0.391 0.392 0.367 0.385 0.450 0.451 0.398 0.404 0.374 0.387 0.380 0.389 0.439 0.450 0.426 0.441
336 0.396 0.402 0.426 0.420 0.424 0.415 0.399 0.410 0.532 0.515 0.428 0.425 0.410 0.411 0.413 0.413 0.490 0.485 0.445 0.459
720 0.462 0.439 0.491 0.459 0.487 0.450 0.454 0.439 0.666 0.589 0.487 0.461 0.478 0.450 0.474 0.453 0.595 0.550 0.543 0.490
Avg 0.382 0.394 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.485 0.481 0.448 0.452
ETTm2 96 0.175 0.258 0.180 0.264 0.182 0.265 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292 0.286 0.377 0.203 0.287
192 0.240 0.302 0.250 0.309 0.246 0.304 0.241 0.302 0.414 0.492 0.290 0.364 0.249 0.309 0.284 0.362 0.399 0.445 0.269 0.328
336 0.306 0.344 0.311 0.348 0.307 0.342 0.305 0.343 0.597 0.542 0.377 0.422 0.321 0.351 0.369 0.427 0.637 0.591 0.325 0.366
720 0.404 0.403 0.412 0.407 0.407 0.398 0.402 0.400 1.730 1.042 0.558 0.524 0.408 0.403 0.554 0.522 0.960 0.735 0.421 0.415
Avg 0.281 0.326 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.571 0.537 0.305 0.349
ETTh1 96 0.382 0.399 0.386 0.405 0.386 0.395 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400 0.654 0.599 0.376 0.419
192 0.437 0.431 0.441 0.436 0.437 0.424 0.460 0.445 0.471 0.474 0.525 0.492 0.436 0.429 0.437 0.432 0.719 0.631 0.420 0.448
336 0.475 0.450 0.487 0.458 0.479 0.446 0.501 0.466 0.570 0.546 0.565 0.515 0.491 0.469 0.481 0.459 0.778 0.659 0.459 0.465
720 0.492 0.474 0.503 0.491 0.481 0.470 0.500 0.488 0.653 0.621 0.594 0.558 0.521 0.500 0.519 0.516 0.836 0.699 0.506 0.507
Avg 0.446 0.438 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.552 0.541 0.507 0.458 0.450 0.456 0.452 0.747 0.647 0.440 0.460
ETTh2 96 0.288 0.340 0.297 0.349 0.288 0.338 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387 0.707 0.621 0.358 0.397
192 0.371 0.396 0.380 0.400 0.374 0.390 0.388 0.400 0.877 0.656 0.528 0.509 0.402 0.414 0.477 0.476 0.860 0.689 0.429 0.439
336 0.412 0.429 0.428 0.432 0.415 0.426 0.426 0.433 1.043 0.731 0.643 0.571 0.452 0.452 0.594 0.541 1.000 0.744 0.496 0.487
720 0.421 0.441 0.427 0.445 0.420 0.440 0.431 0.446 1.104 0.763 0.874 0.679 0.462 0.468 0.831 0.657 1.249 0.838 0.463 0.474
Avg 0.373 0.401 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.954 0.723 0.437 0.449
ECL 96 0.147 0.246 0.148 0.240 0.201 0.281 0.195 0.285 0.219 0.314 0.237 0.329 0.168 0.272 0.197 0.282 0.247 0.345 0.193 0.308
192 0.166 0.261 0.162 0.253 0.201 0.283 0.199 0.289 0.231 0.322 0.236 0.330 0.184 0.289 0.196 0.285 0.257 0.355 0.201 0.315
336 0.184 0.279 0.178 0.269 0.215 0.298 0.215 0.305 0.246 0.337 0.249 0.344 0.198 0.300 0.209 0.301 0.269 0.369 0.214 0.329
720 0.222 0.311 0.225 0.317 0.257 0.331 0.256 0.337 0.280 0.363 0.284 0.373 0.220 0.320 0.245 0.333 0.299 0.390 0.246 0.355
Avg 0.179 0.274 0.178 0.270 0.219 0.298 0.216 0.304 0.244 0.334 0.251 0.344 0.192 0.295 0.212 0.300 0.268 0.365 0.214 0.327
Exchange 96 0.085 0.204 0.086 0.206 0.093 0.217 0.088 0.205 0.256 0.367 0.094 0.218 0.107 0.234 0.088 0.218 0.267 0.396 0.148 0.278
192 0.174 0.298 0.177 0.299 0.184 0.307 0.176 0.299 0.470 0.509 0.184 0.307 0.226 0.344 0.176 0.315 0.351 0.459 0.271 0.315
336 0.317 0.411 0.331 0.417 0.351 0.432 0.301 0.397 1.268 0.883 0.349 0.431 0.367 0.448 0.313 0.427 1.324 0.853 0.460 0.427
720 0.820 0.681 0.847 0.691 0.886 0.714 0.901 0.714 1.767 1.068 0.852 0.698 0.964 0.746 0.839 0.695 1.058 0.797 1.195 0.695
Avg 0.349 0.398 0.360 0.403 0.378 0.417 0.367 0.404 0.940 0.707 0.370 0.413 0.416 0.443 0.354 0.414 0.750 0.626 0.519 0.429
Traffic 96 0.424 0.278 0.395 0.268 0.649 0.389 0.544 0.359 0.522 0.290 0.805 0.493 0.593 0.321 0.650 0.396 0.788 0.499 0.587 0.366
192 0.441 0.289 0.417 0.276 0.601 0.366 0.540 0.354 0.530 0.293 0.756 0.474 0.617 0.336 0.598 0.370 0.789 0.505 0.604 0.373
336 0.461 0.293 0.443 0.283 0.609 0.369 0.551 0.358 0.558 0.305 0.762 0.477 0.629 0.336 0.605 0.373 0.797 0.508 0.621 0.283
720 0.482 0.307 0.467 0.302 0.647 0.387 0.586 0.375 0.589 0.328 0.719 0.449 0.640 0.350 0.645 0.394 0.841 0.523 0.626 0.382
Avg 0.452 0.291 0.428 0.282 0.626 0.378 0.555 0.362 0.550 0.304 0.760 0.473 0.620 0.336 0.625 0.383 0.804 0.509 0.610 0.376
Weather 96 0.157 0.209 0.174 0.214 0.192 0.232 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255 0.221 0.306 0.217 0.296
192 0.206 0.251 0.221 0.254 0.240 0.271 0.225 0.259 0.206 0.277 0.242 0.298 0.219 0.261 0.237 0.296 0.261 0.340 0.276 0.336
336 0.267 0.293 0.278 0.296 0.292 0.307 0.278 0.297 0.272 0.335 0.287 0.335 0.280 0.306 0.283 0.335 0.309 0.378 0.339 0.380
720 0.347 0.346 0.358 0.349 0.364 0.353 0.354 0.348 0.398 0.418 0.351 0.386 0.365 0.359 0.345 0.381 0.377 0.427 0.403 0.428
Avg 0.244 0.274 0.258 0.279 0.272 0.291 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317 0.292 0.363 0.309 0.360
Solar-Energy 96 0.202 0.235 0.203 0.237 0.322 0.339 0.234 0.286 0.310 0.331 0.312 0.399 0.250 0.292 0.290 0.378 0.237 0.344 0.242 0.342
192 0.239 0.262 0.233 0.261 0.359 0.356 0.267 0.310 0.734 0.725 0.339 0.416 0.296 0.318 0.320 0.398 0.280 0.380 0.254 0.272
336 0.251 0.272 0.248 0.273 0.397 0.369 0.290 0.315 0.750 0.735 0.368 0.430 0.319 0.330 0.353 0.415 0.304 0.389 0.282 0.376
720 0.248 0.275 0.249 0.275 0.397 0.356 0.289 0.317 0.769 0.765 0.370 0.425 0.338 0.337 0.356 0.413 0.308 0.388 0.357 0.427
Avg 0.235 0.261 0.233 0.262 0.369 0.270 0.307 0.641 0.6398 0.347 0.417 0.301 0.319 0.330 0.401 0.282 0.375 0.281 0.291 0.381
$\mathrm{1^{st}Count}$ 22 20 11 11 3 11 5 3 0 0 0 0 0 0 0 0 0 0 4 0
Table 2: Ablations of PDMLP. We replace the components of PDMLP one by one and explore the performance of different MLPs. $Avg$ represents the average results across all four forecast lengths. For clarity, the best results are highlighted in bold.
Dateset ECL Solar-Energy Traffic
Model Metric 96 192 336 720 Avg 96 192 336 720 Avg 96 192 336 720 Avg
PDMLP MSE 0.147 0.166 0.184 0.222 0.179 0.202 0.239 0.251 0.248 0.235 0.424 0.441 0.461 0.482 0.452
MAE 0.246 0.261 0.279 0.311 0.274 0.235 0.262 0.272 0.275 0.261 0.278 0.289 0.293 0.307 0.291
PDMLP1 MSE 0.152 0.168 0.192 0.237 0.187 0.217 0.252 0.281 0.274 0.256 0.469 0.484 0.508 0.548 0.502
MAE 0.252 0.265 0.291 0.325 0.283 0.260 0.282 0.300 0.291 0.283 0.324 0.332 0.354 0.365 0.343
PDMLP2 MSE 0.151 0.169 0.187 0.226 0.183 0.210 0.243 0.263 0.261 0.244 0.522 0.525 0.539 0.570 0.539
MAE 0.254 0.268 0.285 0.319 0.281 0.256 0.276 0.290 0.288 0.277 0.342 0.353 0.360 0.451 0.376
PDMLP3 MSE 0.174 0.180 0.201 0.242 0.199 0.222 0.253 0.291 0.289 0.263 0.448 0.459 0.491 0.523 0.480
MAE 0.257 0.263 0.283 0.317 0.280 0.258 0.277 0.300 0.297 0.283 0.290 0.292 0.309 0.327 0.304
DPMLP MSE 0.151 0.168 0.189 0.227 0.183 0.218 0.249 0.263 0.271 0.250 0.480 0.495 0.514 0.554 0.510
MAE 0.251 0.265 0.289 0.320 0.281 0.270 0.292 0.304 0.299 0.291 0.328 0.339 0.346 0.362 0.343
PMLP MSE 0.153 0.170 0.194 0.242 0.189 0.219 0.257 0.274 0.267 0.254 0.474 0.482 0.512 0.557 0.506
MAE 0.254 0.266 0.292 0.330 0.285 0.249 0.288 0.295 0.289 0.280 0.328 0.329 0.342 0.369 0.342
PMLP1 MSE 0.153 0.170 0.197 0.245 0.191 0.225 0.268 0.287 0.285 0.266 0.484 0.491 0.516 0.554 0.510
MAE 0.252 0.268 0.296 0.335 0.287 0.262 0.290 0.303 0.301 0.289 0.334 0.336 0.350 0.370 0.347
PMLP2 MSE 0.153 0.169 0.188 0.227 0.184 0.221 0.266 0.267 0.269 0.255 0.526 0.531 0.528 0.556 0.535
MAE 0.257 0.268 0.288 0.319 0.283 0.248 0.290 0.280 0.283 0.275 0.345 0.350 0.349 0.363 0.351
PMLP3 MSE 0.179 0.185 0.201 0.243 0.202 0.234 0.255 0.290 0.328 0.276 0.455 0.471 0.490 0.517 0.483
MAE 0.260 0.267 0.284 0.317 0.282 0.269 0.278 0.299 0.321 0.291 0.293 0.298 0.308 0.322 0.305
3.4 Projection and Loss Function
We adopt a channel independent approach to implement $\mathrm{Projection}:\mathbb{R}^{d_{model}}\to\mathbb{R}^{T}$ by MLP, where each channel shares parameters.
Our loss function is calculated by the Mean Square Error (MSE) between the model prediction $\hat{x}$ and the ground truth $x$:
$\mathcal{L}=\frac{1}{M}\sum_{M}^{i=1}\left\|\hat{x}^{i}_{L+1:L+T}-x^{i}_{L+1:L+T}\right\|^{2}_{2}$ (2)
and the loss is propagated back from the Projection output across the entire model.
4 Experiments
We comprehensively evaluated the proposed PDMLP in various time series forecasting applications, verifying the generality of the proposed framework, and have further delved into the investigation of the effectiveness of individual components of PDMLP when applied in LSTF tasks.
Datasets: We evaluated the performance of our proposed PDMLP on 8 commonly used LSTF benchmark datasets: Solar Energy Lai et al. (2018), Weather, Traffic, Electricity (ECL) Wu et al. (2021), and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) Zhou et al. (2021).
Baselines and Metrics: We selected nine widely acknowledged SOTA forecasting models for our benchmarking analysis, which comprises the Transformer-based models: iTransformer Liu et al. (2023), PatchTST Nie et al. (2022), Crossformer Zhang & Yan (2022), and FEDformer Zhou et al. (2022); the CNN-based models: TimeNet Wu et al. (2022); and the significant MLP-based models: DLinear Zeng et al. (2023), TiDE Das et al. (2023), and RLinear Li et al. (2023a). To evaluate the performance of these models, we used widely used evaluation metrics: MSE and Mean Absolute Error (MAE).
4.1 Multivariate Long-Term Forecasting
Table  1 presents the results of MTS forecasting, with lower MSE/MAE values indicating greater predictive accuracy. Overall, the model we propose has achieved a surprisingly efficacious outcome, realizing optimal performance across all or the majority of forecast horizons. Out of a total of 90 metrics, a remarkable count of 42 SOTA results were attained, surpassing the second-ranked iTransformer by nearly 50%. This suggests that the Attention mechanism may not necessarily be the optimal choice for LSTF tasks, as even simple linear models are capable of achieving impressive results. Further analysis indicates that channel independence models, represented by PatchTST, often failed to perform to their potential, thus implying the indispensable role of inter-variable interactions in MTS analysis. Notably, despite explicitly capturing multivariate correlations, the performance of Crossformer did not meet expectations, underscoring that inappropriate utilization of inter-variable correlations can negatively impact model efficacy. Hence, we conclude that skillful leverage of interactions between variables is vital and, in certain scenarios, simple linear models may suffice to deliver outstanding performance compared to the more complex Transformer architectures.
4.2 Model Analysis
4.2.1 Ablation Study
As shown in the Table  2, we investigate the impact of each module of PDMLP on performance. DPMLP follows the conventional decomposition approach, which involves decomposing the time series first and then embedding it before separately handling the prediction of the trend and seasonal components. PMLP eliminates the decomposition module, directly inputs the embedded original sequence into the model for prediction. In both PDMLP1 and PMLP1, the MPE is removed, and the embedding method is altered to a single linear layer. PDMLP2 and PMLP2, on the other hand, eliminate the dot product in the MLP among variables and replace it with a simple residual connection, while PDMLP3 and PMLP3 completely remove the MLP among variables, canceling the interaction between them. We only change the corresponding modules and keep all other settings unchanged, with the input length still set to 96.
@@ 4.2.1 Ablation Study
4.2.2 Increasing look-back Window
@@
5 Conclusions

### v3

@@ Introduction
Methods PatchMLP TimeMixer iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear SCINet FEDformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 0.374 0.382 0.381 0.395 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.485 0.481 0.448 0.452
ETTm2 0.269 0.311 0.275 0.323 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.571 0.537 0.305 0.349
ETTh1 0.438 0.429 0.447 0.440 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.552 0.541 0.507 0.458 0.450 0.456 0.452 0.747 0.647 0.440 0.460
ETTh2 0.349 0.378 0.364 0.395 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.954 0.723 0.437 0.449
ECL 0.171 0.265 0.182 0.272 0.178 0.270 0.219 0.298 0.216 0.304 0.244 0.334 0.251 0.344 0.192 0.295 0.212 0.300 0.268 0.365 0.214 0.327
Traffic 0.417 0.273 0.484 0.297 0.428 0.282 0.626 0.378 0.555 0.362 0.550 0.304 0.760 0.473 0.620 0.336 0.625 0.383 0.804 0.509 0.610 0.376
Weather 0.231 0.256 0.240 0.271 0.258 0.279 0.272 0.291 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317 0.292 0.363 0.309 0.360
Solar-Energy 0.211 0.261 0.216 0.280 0.233 0.262 0.369 0.270 0.307 0.641 0.6398 0.347 0.417 0.301 0.319 0.330 0.401 0.282 0.375 0.281 0.291 0.381
$\mathrm{1^{st}Count}$ 8 8 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Table 2: Ablations of PatchMLP. We replace the components of PatchMLP one by one and explore the performance of different MLPs. All the results are averaged from 4 different prediction lengths. A check mark ✓ and a wrong mark ✗ indicate with and without certain components respectively.
Case Decompose MPE Dot Product Intra Variable ECL Traffic Solar-Energy Weather ETTm1
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
1 0.171 0.265 0.417 0.273 0.211 0.277 0.231 0.256 0.374 0.382
2 0.183 0.279 0.431 0.282 0.223 0.288 0.241 0.270 0.383 0.395
3 0.177 0.271 0.426 0.280 0.219 0.283 0.237 0.262 0.383 0.387
4 0.179 0.276 0.426 0.284 0.218 0.287 0.242 0.266 0.382 0.393
5 0.186 0.275 0.429 0.287 0.223 0.288 0.243 0.266 0.384 0.390
6 0.198 0.284 0.442 0.298 0.233 0.301 0.255 0.274 0.396 0.402
7 0.193 0.280 0.435 0.295 0.230 0.293 0.252 0.275 0.390 0.396
8 0.194 0.284 0.439 0.296 0.234 0.301 0.253 0.274 0.395 0.398
9 0.180 0.275 0.429 0.281 0.219 0.287 0.240 0.265 0.382 0.390
Baselines and Metrics: We selected nine widely acknowledged SOTA forecasting models for our benchmarking analysis, which comprises the Transformer-based models: iTransformer (Liu et al. 2023), PatchTST (Nie et al. 2022), Crossformer (Zhang and Yan 2022), and FEDformer (Zhou et al. 2022); the CNN-based models: TimeNet (Wu et al. 2022), SCINet (Liu et al. 2022a); and the significant MLP-based models: Timemixer (Wang et al. 2024), DLinear (Zeng et al. 2023), TiDE (Das et al. 2023), and RLinear (Li et al. 2023a). To evaluate the performance of these models, we used widely used evaluation metrics: MSE and Mean Absolute Error (MAE).
Multivariate Long-Term Forecasting
Table  1 presents the results of MTS forecasting, with lower MSE/MAE values indicating greater predictive accuracy. Overall, the model we propose has achieved a surprisingly efficacious outcome, realizing optimal performance across all datasets. Out of a total of 16 benchmarks, we achieved 100% state-of-the-art (SOTA) results, surpassing all Transformer architectures. This suggests that the Attention mechanism may not necessarily be the optimal choice for LSTF tasks, as even simple linear models are capable of achieving impressive results. Further analysis indicates that channel independence models, represented by PatchTST, often failed to perform to their potential, thus implying the indispensable role of inter-variable interactions in MTS analysis. Notably, despite explicitly capturing multivariate correlations, the performance of Crossformer did not meet expectations, underscoring that inappropriate utilization of inter-variable correlations can negatively impact model efficacy. Hence, we conclude that skillful leverage of interactions between variables is vital and, in certain scenarios, simple linear models may suffice to deliver outstanding performance compared to the more complex Transformer architectures.
Model Analysis
Ablation Study
As shown in the Table  2, we investigate the impact of each module of PatchMLP on performance. 9 follows the conventional decomposition approach, which involves decomposing the time series first and then embedding it before separately handling the prediction of the trend and seasonal components. 5 eliminates the decomposition module, directly inputs the embedded original sequence into the model for prediction. In both 2 and 6, the MPE is removed, and the embedding method is altered to a single linear layer. 3 and 7, on the other hand, eliminate the dot product in the MLP among variables and replace it with a simple residual connection, while 4 and 8 completely remove the MLP among variables, canceling the interaction between them. We only change the corresponding modules and keep all other settings unchanged, with the input length still set to 96.
@@
Increasing look-back Window
@@
Hyperparameter Sensitivity
We evaluated the hypersensitivity of the PatchMLP concerning the following parameters: learning rate (lr), the number of blocks (layers) N in the MLP, and the hidden dimension D of the embedding. The results are illustrated in Figure  6. We observed that the performance of PatchMLP is not particularly sensitive to these hyperparameters, as evidenced by the relatively inconspicuous variations in performance. It is advisable that both the number of blocks and the hidden dimension size not be excessively large, which may be attributable to the increase in the total number of parameters that require learning with larger values of these hyperparameters. However, the hidden dimension size should not be excessively small either, to avoid a disproportionate compression of features into an insufficiently capacious latent vector.
Conclusions